OpenStudy (anonymous):

slope of (3, -4) and (-2, 1)?

OpenStudy (anonymous):

y=-x-1

OpenStudy (anonymous):

use this formula m = y2- y1 / x2 - x1

OpenStudy (anonymous):

yes an then plug it into y-y1=m(x-x1) where m=slope

OpenStudy (anonymous):

then you get your y=mx+b